# ContextGen: Contextual Layout Anchoring for Identity-Consistent Multi-Instance Generation

### v2

@@ Abstract
Multi-instance image generation (MIG) remains a significant challenge for modern diffusion models due to key limitations in achieving precise control over object layout and preserving the identity of multiple distinct subjects. To address these limitations, we introduce ContextGen, a novel Diffusion Transformer framework for multi-instance generation that is guided by both layout and reference images. Our approach integrates two key technical contributions: a Contextual Layout Anchoring (CLA) mechanism that incorporates the composite layout image into the generation context to robustly anchor the objects in their desired positions, and Identity Consistency Attention (ICA), an innovative attention mechanism that leverages contextual reference images to ensure the identity consistency of multiple instances. To address the absence of a large-scale, high-quality dataset for this task, we introduce IMIG-100K, the first dataset to provide detailed layout and identity annotations specifically designed for Multi-Instance Generation. Extensive experiments demonstrate that ContextGen sets a new state-of-the-art, outperforming existing methods especially in layout control and identity fidelity.
@@ 1 Introduction
Diffusion-based models (NEURIPS2020_4c5bcfec) have significantly expanded the horizons of image customization, with many recent systems (e.g., FLUX (flux2024)) adopting the Diffusion Transformer (DiT) (Peebles2022DiT) framework for its enhanced generation quality. Recent developments in subject-driven image generation, such as OmniGen2 (wu2025omnigen2), and layout-to-image synthesis, exemplified by MS-Diffusion (wang2025msdiffusion), have further broadened the scope of customization, enabling control over both content and composition in generated images.
@@
Our method achieves state-of-the-art performance across three benchmarks. On (1) COCO-MIG (zhou2024migc), it improves instance-level success rate by +3.3% and spatial accuracy (mIoU) by +5.9% over prior art. For (2) LayoutSAM-Eval (zhang2024creatilayout), it attains the highest scores in texture and color fidelity, demonstrating superior detail preservation. Most notably, on (3) LAMICBench++ (chen2025lamic), our approach outperforms all open-source models by +1.3% average score and even surpasses commercial systems like GPT-4o in identity retention (+13.3%). These gains validate CLA’s layout robustness and ICA’s effectiveness in multi-instance scenarios.
@@ 2.1 Diffusion Models
Diffusion models have evolved from UNet architectures (NEURIPS2020_4c5bcfec; Rombach_2022_CVPR) to transformer-based approaches like DiT (Peebles2022DiT), enabling scalable multimodal generation as seen in Stable Diffusion 3 (esser2024scalingrectifiedflowtransformers). FLUX (flux2024) further advanced this by unifying visual and textual inputs through multi-modal attention mechanism.
@@ 2.2 Instance-level Controllable Image Generation
GLIGEN (li2023gligen) pioneered the layout-to-image generation paradigm. Follow-up studies utilizing UNet-based methods like InstanceDiffusion (wang2024instancediffusion) and MIGC (migc++), or DiT-based approaches like EliGen (zhang2025eligenentitylevelcontrolledimage) and 3DIS  (zhou20243dis), have demonstrated enhanced capabilities in handling multiple instances. Current state-of-the-art frameworks like OmniGen2 (wu2025omnigen2) and DreamO (mou2025dreamo) process multi-subject conditions via integrated token sequences but face identity degradation with many subjects. While MS-Diffusion (wang2025msdiffusion) and LAMIC (chen2025lamic) combine reference-driven generation with layout control, challenges remain in layout precision and identity consistency.
@@ 3.1 Preliminaries
Recent architectures have replaced modality-specific cross-attention with unified multimodal processing. The MM-Attention operation concatenates image tokens $\mathbf{t}_{\text{image}}$ and text embeddings $\mathbf{t}_{\text{text}}$ into a single sequence $\mathbf{T}=[\mathbf{t}_{\text{text}},\mathbf{t}_{\text{image}}]$, enabling joint self-attention across modalities. Stable Diffusion 3/3.5 (esser2024scalingrectifiedflowtransformers; sd3.5) and FLUX (flux2024), treat all modalities within a shared latent space. The framework naturally supports in-context learning by allowing arbitrary interleaving of visual and textual tokens, while maintaining stable gradient flow across modalities during end-to-end training.
@@
To address the permutation-invariance of the Transformer architecture, Rotatory Position Embedding (RoPE) (su2023roformerenhancedtransformerrotary) was introduced to encode relative positional information. Adapting this for a unified multimodal space, the FLUX.1-Dev architecture proposes a novel extension of RoPE that employs a ternary position encoding scheme. This scheme assigns a position index $\mathbf{p}_{i}=(m,i,j)$ to each token in the sequence. The first component $m$ is set to 0 and is retained for further use. For text tokens, the spatial coordinates $(i,j)$ are fixed at $(0,0)$, while for image tokens, they correspond to the spatial coordinates $(i,j)$ in the 2D noise latent space. This set of position indices $\{\mathbf{p}_{i}\}$ for the sequence forms a position index matrix $\mathbf{P}$.
@@
$\text{MM-Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{(\mathbf{R}\mathbf{Q})(\mathbf{R}\mathbf{K})^{\top}}{\sqrt{d}}\odot\mathbf{M}\right)\mathbf{V}$ (1)
The symbol $\odot$ denotes element-wise multiplication. In the FLUX.1 series, a self-attention mechanism is employed where queries, keys, and values are all derived from the unified token sequence $\mathbf{T}$, and $\mathbf{M}$ is an all-True matrix, enabling full attention across all tokens.
@@ 3.2 Contextual Attention with Layout Anchoring and Identity Preservation
Recent studies in image-to-image (I2I) tasks have demonstrated the effectiveness of using a diptych, a side-by-side reference image pair, to guide diffusion models (shin2024large; song2025insert; zhang2025ICEdit). Building upon this, our framework introduces a novel layout control strategy by integrating a composite layout image into the generation context. This can be done either by a user-defined composition, which offers greater control and is often more aligned with specific user intent, or by our automated sorting algorithm (mentioned in Sec. A.2) based on the occlusion ratio of all instances. This composite diptych serves as the primary input for our Contextual Layout Anchoring (CLA) mechanism, which is designed to enforce a robust spatial structure by anchoring objects to their desired locations.
However, relying solely on this composite layout image presents challenges. As illustrated in Sec. B.5, in scenarios with instance overlap, the process of compositing may result in information loss or detail degradation. To address this, we integrate the original, high-fidelity reference images alongside the diptych. Then the unified token sequence $\mathbf{T}$ mentioned in Sec. 3.1 is constructed as:
$\mathbf{T}=[\mathbf{t}_{\text{text}},\mathbf{t}_{\text{image}},\mathbf{t}_{\text{layout}},\mathbf{t}_{\text{ref}_{1}},\cdots,\mathbf{t}_{\text{ref}_{N}}]$ (2)
@@
Inspired by the functional specialization observed in DiT layers (zhou2025dreamrenderer; zhang2024creatilayout), we propose a hierarchical attention architecture to process the unified token sequence. As shown in the middle panel of Fig. 2, the CLA mechanism operates in the front and back layers, focusing primarily on global context and structural composition. The CLA mask, detailed in the right panel of Fig. 2, ensures broad communication across the text, image, and layout modalities. Using the token sets defined ($\mathcal{T}\!=\!\{\mathbf{t}_{\text{text}}\}$, $\mathcal{I}\!=\!\{\mathbf{t}_{\text{image}}\}$, $\mathcal{L}\!=\!\{\mathbf{t}_{\text{layout}}\}$, and $\mathcal{R}_{n}\!=\!\{\mathbf{t}_{\text{ref}_{n}}\}$) and reference bounding boxes $\{B_{n}\}_{n=1}^{N}$, the attention mask for CLA is defined as:
$\text{M}_{\text{CLA}}(q,k)=\begin{cases}1&\text{if }q\in\mathcal{T}\cup\mathcal{I}\cup\mathcal{L}\text{ and }k\in\mathcal{T}\cup\mathcal{I}\cup\mathcal{L}\\ &\text{or }q\in\mathcal{R}_{n}\text{ and }k\in\mathcal{T}\cup\mathcal{R}_{n}\\ 0&\text{otherwise}\end{cases}$ (3)
where $q$ and $k$ are arbitrary tokens from the query and key sequences, respectively.
@@
While the front and back layers perform global spatial anchoring, we introduce the ICA mechanism in the middle layers to facilitate detailed, instance-level identity injection. As detailed in the right panel of Fig. 2, ICA operates by applying a specialized attention mask, $\text{M}_{\text{ICA}}$, for tokens located within a specific bounding box. For a query token $q\in B_{n}$, the attention mask is defined as:
$\text{M}_{\text{ICA}}(q,k)=\begin{cases}1&\text{if }k\in\mathcal{T}\cup B_{n}\cup\mathcal{R}_{n}\\ 0&\text{otherwise}\\ \end{cases}\qquad\text{if }q\in B_{n}$ (4)
The core function of $\text{M}_{\text{ICA}}$ is the forced connection between $q$ and its corresponding reference tokens $\mathcal{R}_{n}$, ensuring reliable identity transfer. Tokens outside any bounding box (i.e., background) default to the mask used by CLA. This hierarchical strategy effectively transitions our framework from global layout control to refined instance-level identity preservation.
@@
The ternary position encoding scheme described in Sec. 3.1 was extended in FLUX.1-Kontext (labs2025flux1kontextflowmatching) to handle image editing, where the first component of the position index, $m$, was set to 1 for edit tokens. Inspired by this work and other existing work (wu2025less) that shows providing distinct and non-overlapping position indices for each image sequence significantly improves the model’s ability to differentiate between various images, we propose a refined position encoding strategy to systematically structure the relationships within our unified token sequence $\mathbf{T}$ (Eq. 2).
@@ 3.3 IMIG-100K: An Image-Guided Multi-Instance-Generation Dataset
High-fidelity image-guided multi-instance generation is severely limited by the lack of suitable training data. While existing large-scale datasets  (lin2015microsoftcococommonobjects; 5206848) provide diverse instances, they often lack the aesthetic quality and annotation granularity required for modern diffusion models. Conversely, recent subject-driven datasets (tan2025ominicontrol2efficientconditioningdiffusion; xiao2024omnigenunifiedimagegeneration) exhibit high visual quality but are limited by their low instance multiplicity per image. A brief survey on related datasets is provided in Sec. C.1. To bridge this gap, we introduce IMIG-100K, a new large-scale dataset created using the FLUX framework (flux2024). This dataset is specifically designed to support multi-instance generation by providing high-resolution, high-fidelity data with precise layout and reference images.
@@
    Basic Instance Composition (50K samples): This subset focuses on foundational compositional skills. The ground truth images are generated by the text-to-image model FLUX.1-Dev  (flux2024), and we derive reference images using detection and segmentation models (liu2023grounding; ravi2024sam2; DeRIS). These reference images undergo minimal post-processing, including basic lighting adjustments.
@@
    Flexible Composition with References (10K samples): Unlike the previous two subsets, this unique subset is designed to train the model’s robustness in handling low-consistency inputs. We first generate individual reference instances using the FLUX.1-Dev model. These are then composited into ground truth scenes by subject-driven models (wu2025less; mou2025dreamo), allowing for a much greater degree of flexibility and transformation in the composited instances relative to their original references. A key step involves rigorous filtering to ensure identity consistency from the references (guo2021sample; oquab2023dinov2).
All textual prompts are generated by advanced large language models (deepseekai2025deepseekv3technicalreport; comanici2025gemini25pushingfrontier; openai2024gpt4ocard), ensuring diverse and high-quality descriptions. Further details on the dataset are provided in Sec. C.2.
@@ 4.1 Experimental Setting
We initialize the model with FLUX.1-Kontext (labs2025flux1kontextflowmatching) without introducing additional parameters and fine-tune it using LoRA (Low-Rank Adaptation) (hu2021loralowrankadaptationlarge) with LoRA Rank 512. We perform training on 4× NVIDIA A100 GPUs with a total batch size of 16. The model is tuned on the three hierarchical sub-datasets described in Sec. 3.3 for 5K steps, employing the Prodigy optimizer (mishchenko2024prodigy) with its default learning rate. We also employ Direct Preference Optimization (DPO) (rafailov2024directpreferenceoptimizationlanguage) (detailed in Sec. 4.4) to refine text-visual alignment and user preference. These enable the model to evolve from mastering simple compositions to synthesizing complex multi-instance scenes.
@@
(1) LAMICBench++: A specialized benchmark for evaluating identity preservation and feature consistency in subject-driven generation. We extend the multi-image composition benchmark from LAMICBench (chen2025lamic), aggregating multi-category reference images (humans, animals, objects, etc.) from established datasets including XVerseBench (chen2025xverse), DreamBench++(peng2024dreambench) and MS-Bench (wang2025msdiffusion). In particular, we construct a dataset of 160 cases in total, including 50 cases with 2 reference images, 40 with 3, 30 with 4, 20 with 5, and 20 with over 5 reference images. These cases are divided into two categories: Fewer Subjects ($\leq 3$ reference images) and More Subjects ($\geq 4$ reference images). In this benchmark, we adapt and slightly modified the four evaluation metrics from the original work: (1) Global text-image consistency (ITC) evaluated through visual-question-answering (VQA)  (ye2024mplugowl3longimagesequenceunderstanding), with approximately 2K questions (4-12 per item); (2) Object preservation (IPS) (liu2023grounding; oquab2023dinov2); (3) Facial identity retention (IDS) (guo2021sample); (4) Aesthetic quality (AES) (schuhmann2023aes).
(2) COCO-MIG (zhou2024migc): A benchmark designed to evaluate spatial and attribute accuracy in layout-to-image generation, comprising 800 images from COCO Dataset (lin2015microsoftcococommonobjects) with color-annotated instances. The evaluation metrics include: (1) Global and instance level success rate (SR and I-SR) determined by spatial accuracy (mIoU) and color correctness; (2) Multi-scale semantic consistency through global and local CLIP Scores (G-C and L-C).
(3) LayoutSAM-Eval (zhang2024creatilayout): A open-set benchmark for layout-to-image evaluation, featuring 5K prompts with exhaustive entity-level annotations, from which we filter 1K samples with sufficiently large bounding boxes for reliable instance evaluation. We adapt the original work’s metrics: (1) Fine-grained entity accuracy (spatial, color, textural, shape) evaluated using MLLM (yao2024minicpm); (2) Holistic quality metrics: CLIP Score for semantic alignment and Pick Score (kirstain2023pickapicopendatasetuser) for human preference.

### v3

@@ Abstract
Multi-instance image generation (MIG) remains a significant challenge for modern diffusion models due to key limitations in achieving precise control over object layout and preserving the identity of multiple distinct subjects. To address these limitations, we introduce ContextGen, a novel Diffusion Transformer framework for multi-instance generation that is guided by both layout and reference images. Our approach integrates two key technical contributions: a Contextual Layout Anchoring (CLA) mechanism that incorporates the composite layout image into the generation context to robustly anchor the objects in their desired positions, and Identity Consistency Attention (ICA), an innovative attention mechanism that leverages contextual reference images to ensure the identity consistency of multiple instances. To address the absence of a large-scale, high-quality dataset for this task, we introduce IMIG-100K, the first dataset to provide detailed layout and identity annotations specifically designed for Multi-Instance Generation. Extensive experiments demonstrate that ContextGen sets a new state-of-the-art, outperforming existing methods especially in layout control and identity fidelity. Our code is available at https://github.com/nenhang/ContextGen.
@@ 1 Introduction
Diffusion-based models (Ho et al., 2020) have significantly expanded the horizons of image customization, with many recent systems (e.g., FLUX (Labs, 2024b)) adopting the Diffusion Transformer (DiT) (Peebles and Xie, 2022) framework for its enhanced generation quality. Recent developments in subject-driven image generation, such as OmniGen2 (Wu et al., 2025b), and layout-to-image synthesis, exemplified by MS-Diffusion (Wang et al., 2025), have further broadened the scope of customization, enabling control over both content and composition in generated images.
@@
Our method achieves state-of-the-art performance across three benchmarks. On (1) COCO-MIG (Zhou et al., 2024b), it improves instance-level success rate by +3.3% and spatial accuracy (mIoU) by +5.9% over prior art. For (2) LayoutSAM-Eval (Zhang et al., 2024), it attains the highest scores in texture and color fidelity, demonstrating superior detail preservation. Most notably, on (3) LAMICBench++ (Chen et al., 2025b), our approach outperforms all open-source models by +1.3% average score and even surpasses commercial systems like GPT-4o in identity retention (+13.3%). These gains validate CLA’s layout robustness and ICA’s effectiveness in multi-instance scenarios.
@@ 2.1 Diffusion Models
Diffusion models have evolved from UNet architectures (Ho et al., 2020; Rombach et al., 2022) to transformer-based approaches like DiT (Peebles and Xie, 2022), enabling scalable multimodal generation as seen in Stable Diffusion 3 (Esser et al., 2024). FLUX (Labs, 2024b) further advanced this by unifying visual and textual inputs through multi-modal attention mechanism.
@@ 2.2 Instance-level Controllable Image Generation
GLIGEN (Li et al., 2023) pioneered the layout-to-image generation paradigm. Follow-up studies utilizing UNet-based methods like InstanceDiffusion (Wang et al., 2024) and MIGC (Zhou et al., 2024a), or DiT-based approaches like EliGen (Zhang et al., 2025a) and 3DIS  (Zhou et al., 2024c), have demonstrated enhanced capabilities in handling multiple instances. Current state-of-the-art frameworks like OmniGen2 (Wu et al., 2025b) and DreamO (Mou et al., 2025) process multi-subject conditions via integrated token sequences but face identity degradation with many subjects. While MS-Diffusion (Wang et al., 2025) and LAMIC (Chen et al., 2025b) combine reference-driven generation with layout control, challenges remain in layout precision and identity consistency.
@@ 3.1 Preliminaries
Recent architectures have replaced modality-specific cross-attention with unified multimodal processing. The MM-Attention operation concatenates image tokens $\mathbf{t}_{\text{image}}$ and text embeddings $\mathbf{t}_{\text{text}}$ into a single sequence $\mathbf{T}=[\mathbf{t}_{\text{text}},\mathbf{t}_{\text{image}}]$, enabling joint self-attention across modalities. Stable Diffusion 3/3.5 (Esser et al., 2024; stability.ai, 2024) and FLUX (Labs, 2024b), treat all modalities within a shared latent space. The framework naturally supports in-context learning by allowing arbitrary interleaving of visual and textual tokens, while maintaining stable gradient flow across modalities during end-to-end training.
@@
To address the permutation-invariance of the Transformer architecture, Rotatory Position Embedding (RoPE) (Su et al., 2023) was introduced to encode relative positional information. Adapting this for a unified multimodal space, the FLUX.1-Dev architecture proposes a novel extension of RoPE that employs a ternary position encoding scheme. This scheme assigns a position index $\mathbf{p}_{i}=(m,i,j)$ to each token in the sequence. The first component $m$ is set to 0 and is retained for further use. For text tokens, the spatial coordinates $(i,j)$ are fixed at $(0,0)$, while for image tokens, they correspond to the spatial coordinates $(i,j)$ in the 2D noise latent space. This set of position indices $\{\mathbf{p}_{i}\}$ for the sequence forms a position index matrix $\mathbf{P}$.
@@
$\text{MM-Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{(\mathbf{R}\mathbf{Q})(\mathbf{R}\mathbf{K})^{\top}}{\sqrt{d}}\odot\mathbf{M}\right)\mathbf{V},$ (1)
where $\odot$ denotes element-wise multiplication. In the FLUX.1 series, a self-attention mechanism is employed where queries, keys, and values are all derived from the unified token sequence $\mathbf{T}$, and $\mathbf{M}$ is an all-True matrix, enabling full attention across all tokens.
@@ 3.2 Contextual Attention with Layout Anchoring and Identity Preservation
Recent studies in image-to-image (I2I) tasks have demonstrated the effectiveness of using a diptych, a side-by-side reference image pair, to guide diffusion models (Shin et al., 2024; Song et al., 2025; Zhang et al., 2025c). Building upon this, our framework introduces a novel layout control strategy by integrating a composite layout image into the generation context. This can be done either by a user-defined composition, which offers greater control and is often more aligned with specific user intent, or by our automated sorting algorithm (mentioned in Sec. A.2) based on the occlusion ratio of all instances. This composite diptych serves as the primary input for our Contextual Layout Anchoring (CLA) mechanism, which is designed to enforce a robust spatial structure by anchoring objects to their desired locations.
However, relying solely on this composite layout image presents challenges. As illustrated in Sec. B.5, in scenarios with instance overlap, the process of compositing may result in information loss or detail degradation. To address this, we integrate the original, high-fidelity reference images alongside the diptych. The unified token sequence $\mathbf{T}$ mentioned in Sec. 3.1 is constructed as:
$\mathbf{T}=[\mathbf{t}_{\text{text}},\mathbf{t}_{\text{image}},\mathbf{t}_{\text{layout}},\mathbf{t}_{\text{ref}_{1}},\cdots,\mathbf{t}_{\text{ref}_{N}}].$ (2)
@@
Inspired by the functional specialization observed in DiT layers (Zhou et al., 2025b; Zhang et al., 2024), we propose a hierarchical attention architecture to process the unified token sequence. As shown in the middle panel of Fig. 2, the CLA mechanism operates in the front and back layers, focusing primarily on global context and structural composition. The CLA mask, detailed in the right panel of Fig. 2, ensures broad communication across the text, image, and layout modalities. Using the token sets defined ($\mathcal{T}\!=\!\{\mathbf{t}_{\text{text}}\}$, $\mathcal{I}\!=\!\{\mathbf{t}_{\text{image}}\}$, $\mathcal{L}\!=\!\{\mathbf{t}_{\text{layout}}\}$, and $\mathcal{R}_{n}\!=\!\{\mathbf{t}_{\text{ref}_{n}}\}$) and reference bounding boxes $\{B_{n}\}_{n=1}^{N}$, the attention mask for CLA is defined as:
$\mathbf{M}_{\text{CLA}}(q,k)=\mathbbm{1}\left[(q,k)\in(\mathcal{T}\cup\mathcal{I}\cup\mathcal{L})^{2}\cup\bigcup\limits_{n=1}^{N}(\mathcal{R}_{n}\times(\mathcal{T}\cup\mathcal{R}_{n}))\right],$ (3)
where $q$ and $k$ are arbitrary tokens from the query and key sequences respectively, and $\mathbbm{1}\left[\cdot\right]$ denotes the indicator function.
@@
While the front and back layers perform global spatial anchoring, we introduce the ICA mechanism in the middle layers to facilitate detailed, instance-level identity injection. As detailed in the right panel of Fig. 2, ICA operates by applying a specialized attention mask, $\mathbf{M}_{\text{ICA}}$, for tokens located within a specific bounding box. For a query token $q\in B_{n}$, the attention mask is defined as:
$\mathbf{M}_{\text{ICA}}(q,k)=\mathbbm{1}\left[(q,k)\in\bigcup\limits_{n=1}^{N}(B_{n}\times(\mathcal{T}\cup B_{n}\cup\mathcal{R}_{n}))\cup\{(q,k)\in\mathbf{M}_{\text{CLA}}\mid q\notin\bigcup\limits_{n=1}^{N}B_{n}\}\right].$ (4)
The core function of $\mathbf{M}_{\text{ICA}}$ is the forced connection between $q$ and its corresponding reference tokens $\mathcal{R}_{n}$, ensuring reliable identity transfer. Tokens outside any bounding box (i.e., background) default to the mask used by CLA. This hierarchical strategy effectively transitions our framework from global layout control to refined instance-level identity preservation.
@@
The ternary position encoding scheme described in Sec. 3.1 was extended in FLUX.1-Kontext (Labs et al., 2025) to handle image editing, where the first component of the position index, $m$, was set to 1 for edit tokens. Inspired by this work and other existing work (Wu et al., 2025c) that shows providing distinct and non-overlapping position indices for each image sequence significantly improves the model’s ability to differentiate between various images, we propose a refined position encoding strategy to systematically structure the relationships within our unified token sequence $\mathbf{T}$ (Eq. 2).
@@ 3.3 IMIG-100K: An Image-Guided Multi-Instance-Generation Dataset
High-fidelity image-guided multi-instance generation is severely limited by the lack of suitable training data. While existing large-scale datasets  (Lin et al., 2015; Deng et al., 2009) provide diverse instances, they often lack the aesthetic quality and annotation granularity required for modern diffusion models. Conversely, recent subject-driven datasets (Tan et al., 2025; Xiao et al., 2024a) exhibit high visual quality but are limited by their low instance multiplicity per image. A brief survey on related datasets is provided in Sec. C.1. To bridge this gap, we introduce IMIG-100K, a large-scale dataset created using the FLUX framework (Labs, 2024b). This dataset is specifically designed to support multi-instance generation by providing high-resolution, high-fidelity data with precise layout and reference images.
@@
    Basic Instance Composition (50K samples): This subset focuses on foundational compositional skills. The ground truth images are generated by the text-to-image model FLUX.1-Dev  (Labs, 2024b), and we derive reference images using detection and segmentation models (Liu et al., 2023; Ravi et al., 2024; Dai et al., 2025). These reference images undergo minimal post-processing, including basic lighting adjustments.
@@
    Flexible Composition with References (10K samples): Unlike the previous two subsets, this unique subset is designed to train the model’s robustness in handling low-consistency inputs. We first generate individual reference instances using the FLUX.1-Dev model. These are then composited into ground truth scenes by subject-driven models (Wu et al., 2025c; Mou et al., 2025), allowing for a much greater degree of flexibility and transformation in the composited instances relative to their original references. A key step involves rigorous filtering to ensure identity consistency from the references (Guo et al., 2021; Oquab et al., 2023).
All textual prompts are generated by advanced large language models (DeepSeek-AI, 2025; Comanici et al., 2025; OpenAI, 2024), ensuring diverse and high-quality descriptions. Further details on the dataset are provided in Sec. C.2.
@@ 4.1 Experimental Setting
We initialize the model with FLUX.1-Kontext (Labs et al., 2025) without introducing additional parameters and fine-tune it using LoRA (Low-Rank Adaptation) (Hu et al., 2021) with LoRA Rank 512. We perform training on 4× NVIDIA A100 GPUs with a total batch size of 16. The model is tuned on the three hierarchical sub-datasets described in Sec. 3.3 for 5K steps, employing the Prodigy optimizer (Mishchenko and Defazio, 2024) with its default learning rate. We also employ Direct Preference Optimization (DPO) (Rafailov et al., 2024) (detailed in Sec. 4.4) to refine text-visual alignment and user preference. These enable the model to evolve from mastering simple compositions to synthesizing complex multi-instance scenes.
@@
(1) LAMICBench++: A specialized benchmark for evaluating identity preservation and feature consistency in subject-driven generation. We extend the multi-image composition benchmark from LAMICBench (Chen et al., 2025b), aggregating multi-category reference images (humans, animals, objects, etc.) from established datasets including XVerseBench (Chen et al., 2025a), DreamBench++(Peng et al., 2025) and MS-Bench (Wang et al., 2025). In particular, we construct a dataset of 160 cases in total, including 50 cases with 2 reference images, 40 with 3, 30 with 4, 20 with 5, and 20 with over 5 reference images. These cases are divided into two categories: Fewer Subjects ($\leq 3$ reference images) and More Subjects ($\geq 4$ reference images). In this benchmark, we adapt and slightly modified the four evaluation metrics from the original work: (1) Global text-image consistency (ITC) evaluated through visual-question-answering (VQA)  (Ye et al., 2024), with approximately 2K questions (4-12 per item); (2) Object preservation (IPS) (Liu et al., 2023; Oquab et al., 2023); (3) Facial identity retention (IDS) (Guo et al., 2021); (4) Aesthetic quality (AES) (Schuhmann, 2023).
(2) COCO-MIG (Zhou et al., 2024b): A benchmark designed to evaluate spatial and attribute accuracy in layout-to-image generation, comprising 800 images from COCO Dataset (Lin et al., 2015) with color-annotated instances. The evaluation metrics include: (1) Global and instance level success rate (SR and I-SR) determined by spatial accuracy (mIoU) and color correctness; (2) Multi-scale semantic consistency through global and local CLIP Scores (G-C and L-C).
(3) LayoutSAM-Eval (Zhang et al., 2024): An open-set benchmark for layout-to-image evaluation, featuring 5K prompts with exhaustive entity-level annotations, from which we filter 1K samples with sufficiently large bounding boxes for reliable instance evaluation. We adapt the original work’s metrics: (1) Fine-grained entity accuracy (spatial, color, textural, shape) evaluated using MLLM (Yao et al., 2024); (2) Holistic quality metrics: CLIP Score for semantic alignment and Pick Score (Kirstain et al., 2023) for human preference.